OpenStudy (anonymous):

If it requires 7.0J of work to stretch a particular spring by 1.8cm from its equilibrium length, how much more work will be required to stretch it an additional 4.2cm ?

OpenStudy (abhisar):

Hello @zfower3317 ! \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

Have you tried it yourself ?

OpenStudy (abhisar):

Work done on the spring gets stored as its potential energy. Potential energy of a spring is given by \(\large\sf \frac{1}{2}kx^2\). -->\(\large\sf \frac{1}{2}k(1.8\times10^{-2})^2=7\). Now find the value of k

OpenStudy (abhisar):

Substitute this value in the equation \(\large\sf \frac{1}{2}k(4.2\times10^{-2})^2\). This will give you the work required in displacing the spring by 4.2 cm. Subtract 7 from it and it will give you the answer.

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

yeah i think i get now thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)